OpenStudy (anonymous):

a married couple borrows $50000 in order to purchase a house. they agree to repay the loan in equal quarterly installments for 12 years. that is the amount of each installment, if interest is charges at the rate of 10% per annum on any money owing?

OpenStudy (tkhunny):

This is your basic present value problem \[50,000 = Payment\cdot\frac{1-v^{12*4}}{i}\] where i = 0.10 / 4 and v = 1/(1+i)